# 1]Meta Superintelligence Labs, Menlo Park, CA 94025, USA

### v1

@@ 3.1 Interpreting GPA as Smoothed DiLoCo
As seen in Figure 1(a), increasing the number of inner steps leads to improved performance for single-worker DiLoCo. However, the underlying reasons for this behavior are not understood. By examining DiLoCo from the lens of GPA in equation 8 and comparing it with the more restrictive Nesterov formulation in equation 4, we can develop a deeper intuition for DiLoCo’s inner workings.
Suppose that we increase the number of inner steps in DiLoCo and want to maintain the same level of smoothing on the average iterate $x^{(t)}$. One may attempt to increase $\mu$ in Nesterov (equation 4) to decrease the weight on the current iterate $z^{(t+1)}$. However, since $\mu$ controls both the amount of smoothing in $x^{(t)}$ and the amount of interpolation used to update $y^{(t)}$, strictly increasing $\mu$ would decrease the recency of information from $z^{(t)}$ in $y^{(t)}$ by a factor of $\mu^{2}$, resulting in significantly different algorithmic behavior. Numerically, we validate that tuning $\mu$ alone in Nesterov’s primal averaging formulation is not sufficient to reach the performance of DiLoCo; see Appendix 11.
@@
Tuning GPA from DiLoCo. This intuition provides practical guidelines for converting a tuning for DiLoCo to GPA. Given an optimal number of inner steps $H$ and momentum parameter $\mu$ in DiLoCo, we observe for GPA that $x^{(t+H)}=\mu_{x}^{H}x^{(t)}+(1-\mu_{x})\sum_{k=0}^{H-1}\mu_{x}^{k}z^{(t+H-k)}$. Therefore, to match the coefficient in front of $x^{(t)}$ with DiLoCo, one can set $\mu_{x}=\mu^{1/H}$ while keeping $\mu_{y}\approx\mu$. With commonly used values $\mu=0.9$ and $H=32$, we obtain $\mu_{x}\approx 0.9967$ and $\mu_{y}\approx 0.9$. We leverage this heuristic to determine an effective number of inner steps used in Figure 1. See Table 4 for exact values mapping inner steps in DiLoCo to GPA coefficient $\mu_{x}$.
Tradeoffs with DiLoCo. GPA not only outperforms DiLoCo, but does so with fewer hyperparameters and lower memory requirements. While DiLoCo requires four hyperparameters, e.g., the inner and outer learning rate, momentum hyperparameter, and number of inner steps, GPA reduces this to just three: the learning rate and two momentum parameters. This simplification is possible because DiLoCo’s practical performance is governed by an effective learning rate that couples the effect of the inner and outer learning rates ($\gamma^{(t)}$ and $\tilde{\gamma}$). On the other hand, GPA requires more FLOPs per-iteration, while DiLoCo amortizes its additional compute cost across multiple inner steps.
Table 1: Summary of Nesterov, Schedule-Free, and GPA formulations. Here, $\mu$ is the momentum parameter, $\gamma^{(t)}$ is the learning rate, and $g(\cdot;\xi^{(t)})$ denotes the stochastic gradient.
Formulation Update Equations Notes
Sutskever (Classical) (pmlr-v28-sutskever13) $\begin{aligned} b^{(t)}&=\mu b^{(t-1)}-\gamma^{(t)}g(x^{(t)}+\mu b^{(t-1)};\xi^{(t)})\\ x^{(t+1)}&=x^{(t)}+b^{(t)}\end{aligned}$ Gradient is evaluated at the lookahead point.
Modern (PyTorch/JAX) $\begin{aligned} b^{(t)}&=\mu b^{(t-1)}+g(x^{(t)};\xi^{(t)})\\ x^{(t+1)}&=x^{(t)}-\gamma^{(t)}[\mu b^{(t)}+g(x^{(t)};\xi^{(t)})]\end{aligned}$ Gradient is evaluated at the current point.
Primal Averaging Variant of Nesterov (lanaccel) $\begin{aligned} y^{(t)}&=\mu x^{(t)}+(1-\mu)z^{(t)}\\ z^{(t+1)}&=z^{(t)}-\gamma^{(t)}g(y^{(t)};\xi^{(t)})\\ x^{(t+1)}&=\mu x^{(t)}+(1-\mu)z^{(t+1)}\end{aligned}$ Explicit separation of gradient and model evaluation sequences.
Schedule-Free (schedule-free) $\begin{aligned} y^{(t)}&=\mu x^{(t)}+(1-\mu)z^{(t)}\\ z^{(t+1)}&=z^{(t)}-\gamma g(y^{(t)};\xi^{(t)})\\ x^{(t+1)}&=\frac{t}{t+1}x^{(t)}+\left(1-\frac{t}{t+1}\right)z^{(t+1)}\end{aligned}$ Uniform averaging; learning rate schedule-free.
GPA (Ours) $\begin{aligned} y^{(t)}&=\mu_{y}x^{(t)}+(1-\mu_{y})z^{(t)}\\ z^{(t+1)}&=z^{(t)}-\gamma^{(t)}g(y^{(t)};\xi^{(t)})\\ x^{(t+1)}&=\mu_{x}x^{(t)}+(1-\mu_{x})z^{(t+1)}\end{aligned}$ Decoupled interpolation of gradient and model evaluation sequences; requires learning rate schedule.
@@ 4 Experiments
In this section, we assess the effectiveness of GPA on both language model pre-training and computer vision workloads. For language modeling, we compare against baselines AdamW and DiLoCo, while for computer vision experiments we compare GPA against AdamW. For both DiLoCo and GPA, we use AdamW as the base optimizer (DiLoCo-AdamW and GPA-AdamW, respectively).
4.1 Language Model Pre-Training
We conduct experiments on two scales of Llama models: (1) 160 million parameters and (2) 1 billion parameters. These are pre-trained on the C4 dataset from scratch (2019t5) using a token budget of roughly 3.2 billion and 50 billion tokens, respectively (hoffmann2022training). All of our small experiments are conducted on a single machine equipped with eight H100 GPUs (97 GB of memory) while the large scale model experiments utilize two nodes (with a total of 16 GPUs). Comprehensive details on batch size, sequence length, and hyperparameter sweeps can be found in Appendix 11. Note that the Llama-1B experiments are performed in an overtrained setting.
@@ 4.1 Language Model Pre-Training
Method $H=8$ $H=16$ $H=32$ $H=64$ $H=16$ $H=32$ $H=64$ $H=128$
AdamW 3.2691 3.2691 3.2691 3.2691 2.6886 2.6886 2.6886 2.6886
DiLoCo-AdamW 3.2685 3.2619 3.2836 3.2976 2.6835 2.6765 2.6755 2.6743
GPA-AdamW 3.2672 3.2546 3.2489 3.2664 2.6828 2.6722 2.6619 2.6734
Performance across number of inner steps. In Table 2, we provide the final validation loss for each method for different effective number of inner steps. Consistent with Figure 1(a), GPA-AdamW outperforms both DiLoCo-AdamW and AdamW, except when the number of inner steps is 1. Both DiLoCo-AdamW and GPA-AdamW display U-shaped behavior with respect to the number of inner steps.
@@
Convergence behavior. Figure 3(a) shows the validation loss curves on Llama-160M for AdamW, DiLoCo-AdamW, and GPA-AdamW for the case where the number of inner steps is 32. In this case, $\mu_{x}$ has been tuned to match the number of inner steps; see Table 4 in Appendix 11 for details. In Figure 3(b), we compare the GPA against the baselines (also including ScheduleFree-AdamW) by choosing the best performing runs over all hyperparameter choices including the effective number of inner steps. GPA-AdamW converges faster than both DiLoCo and AdamW throughout the entire training run. The training curves for GPA-AdamW are also noticeably smoother and more stable compared to the other methods. Our hyperparameter sweeps reveal that GPA-AdamW can handle higher learning rates compared to DiLoCo and AdamW, e.g., $1\cdot 10^{-2}$.
@@ 4.2 Vision Transformer Model Training
To validate our method on a computer vision task, we train a ViT-S/16 model from timm on ImageNet with data augmentations from the repository. We train this under two batch size settings: (1) a small-batch setup with a batch size of 4,096 for 300 epochs; and (2) a large-batch setup with a batch size of 16,384 for 200 epochs. We tuned the methods separately in both settings, using the average over 2 random seeds to select the best hyperparameters, then run the best-performing selection on 8 random seeds in total. For all methods, we used gradient clipping with norm 1 and a learning rate schedule consisting of linear warmup over the first 5 epochs, then cosine decay to $\times 0.002$ of the peak learning rate.
Our evaluation in both small-batch and large-batch settings indicate that GPA outperforms AdamW by a clear margin in terms of validation accuracy throughout the course of training (see Figures 4 and 5). For further details on our hyperparameter tuning, see Appendix 11.
Figure 4: Comparison of AdamW and GPA on ImageNet ViT-S/16 from timm with data augmentations using a batch size of 4,096 samples. The optimal configuration for both AdamW and GPA use a learning rate of 0.005 and weight decay of 0.1.
Figure 5: Comparison of AdamW and GPA on ImageNet ViT-S/16 from timm with data augmentations using a batch size of 16,384 samples.
@@ 6 Conclusion
GPA introduces independent interpolation constants for the gradient computation and model evaluation sequences that yield a flexible generalization of Nesterov momentum. On both small and large-scale dense language models, this flexibility allows GPA to outperform single-worker DiLoCo, while removing the complexity of its two-loop structure, simplifying its hyperparameter tuning and reducing its memory requirements.
@@
Finally, GPA’s decoupling of parameters also enables new avenues for distributed training. In DiLoCo, the number of inner steps serves as a coupled hyperparameter for both Lookahead with Nesterov and local SGD, leading to the undesirable finding that increasing the number of inner steps can improve convergence – contrary to standard local SGD intuition. GPA introduces a tunable, continuous smoothing parameter that is independent of the number of local SGD steps, laying a new foundation for re-designing DiLoCo for cross-regional training.

### v3

@@ 3.1 Interpreting GPA as Smoothed DiLoCo
As seen in Figure 1(a), increasing the number of inner steps leads to improved performance for single-worker DiLoCo. However, the underlying reasons for this behavior are not understood. By examining DiLoCo from the lens of GPA in Equation˜8 and comparing it with the more restrictive Nesterov formulation in Equation˜4, we can develop a deeper intuition for DiLoCo’s inner workings.
Suppose that we increase the number of inner steps in DiLoCo and want to maintain the same level of smoothing on the average iterate $x^{(t)}$. One may attempt to increase $\mu$ in Nesterov (Equation˜4) to decrease the weight on the current iterate $z^{(t+1)}$. However, since $\mu$ controls both the amount of smoothing in $x^{(t)}$ and the amount of interpolation used to update $y^{(t)}$, strictly increasing $\mu$ would decrease the recency of information from $z^{(t)}$ in $y^{(t)}$ by a factor of $\mu^{2}$, resulting in significantly different algorithmic behavior. Numerically, we validate that tuning $\mu$ alone in Nesterov’s primal averaging formulation is not sufficient to reach the performance of DiLoCo; see Appendix 11.
@@
Tuning GPA from DiLoCo. This intuition provides practical guidelines for converting a hyperparameter recipe for DiLoCo to GPA. Given an optimal number of inner steps $H$ and momentum parameter $\mu$ in DiLoCo, we observe for GPA that $x^{(t+H)}=\mu_{x}^{H}x^{(t)}+(1-\mu_{x})\sum_{k=0}^{H-1}\mu_{x}^{k}z^{(t+H-k)}$. Therefore, to match the coefficient in front of $x^{(t)}$ with DiLoCo, one can set $\mu_{x}=\mu^{1/H}$ while keeping $\mu_{y}\approx\mu$. With commonly used values $\mu=0.9$ and $H=32$, we obtain $\mu_{x}\approx 0.9967$ and $\mu_{y}\approx 0.9$. We leverage this heuristic to determine an effective number of inner steps used in Figure 1. See Table 6 for exact values mapping inner steps in DiLoCo to GPA coefficient $\mu_{x}$.
Tradeoffs with DiLoCo. GPA not only outperforms DiLoCo, but does so with fewer hyperparameters and lower memory requirements. While DiLoCo requires four hyperparameters, e.g., the inner and outer learning rate, momentum hyperparameter, and number of inner steps, the memory-efficient implementation of GPA reduces this to just three: the learning rate and two momentum parameters. This simplification is possible because DiLoCo’s practical performance is governed by an effective learning rate that couples the effect of the inner and outer learning rates ($\gamma^{(t)}$ and $\tilde{\gamma}$). On the other hand, GPA requires more FLOPs per-iteration, while DiLoCo amortizes its additional compute cost across multiple inner steps (see Appendix 9.3 for additional details).
@@ 4 Experiments
In this section, we assess the effectiveness of GPA on both language model pre-training and computer vision workloads. For language modeling, we compare against baselines AdamW, DiLoCo and Schedule-Free (SF), while for computer vision experiments, we compare GPA against AdamW. For DiLoCo, Schedule-Free and GPA methods, we use AdamW as the base optimizer (DiLoCo-AdamW, Schedule-Free and GPA-AdamW, respectively).
@@
Method $H=8$ $H=16$ $H=32$ $H=64$ $H=128$ $H=16$ $H=32$ $H=64$ $H=128$
[…]
@@
4.1 Language Model Pre-Training
We conduct experiments on multiple scales of Llama models: 160 million and 1 billion parameters. These 160M and 1B models are pre-trained on the C4 dataset from scratch (2019t5) using a token budget of roughly 9.6 billion and 50 billion tokens, respectively (hoffmann2022training). All of our small experiments are conducted on two nodes equipped with two GB200 GPUs (184 GB of memory) while the larger scale model experiments utilize four nodes (with a total of 8 GPUs). Comprehensive details on batch size, sequence length, and hyperparameter sweeps can be found in Appendix 11. Note that both the Llama-160M and 1B experiments are performed in an overtrained setting.
Performance across number of inner steps. In Table 2, we provide the final validation loss for each method for different effective number of inner steps. Consistent with Figure 1(a), GPA-AdamW outperforms both DiLoCo-AdamW and AdamW across all settings of the effective number of inner steps.
Convergence behavior. Figure 3(a) shows the validation loss curves on Llama-160M for AdamW, DiLoCo-AdamW, and GPA-AdamW for the case where the number of inner steps is 32. In this case, $\mu_{x}$ has been tuned to match the number of inner steps; see Table 6 in Appendix 11 for details. In Figure 3(b), we compare GPA against the baselines (also including Schedule-Free-AdamW) by choosing the best performing runs over all hyperparameter choices including the effective number of inner steps. GPA-AdamW converges faster than both DiLoCo and AdamW throughout the entire training run. The training curves for GPA-AdamW are also noticeably smoother and more stable compared to the other methods. We observe that GPA-AdamW can handle higher learning rates compared to DiLoCo.
4.2 Code Generation Model Pre-Training
We also evaluate GPA on an 8 billion parameter Llama model designed for code generation; see Appendix 11.7 for more details. We find that GPA with any $\mu_{y}\in\{0.7,0.8,0.8\}$ outperforms AdamW throughout the course of training and achieves a better final validation loss (see Figure 4, which plots GPA with $\mu_{y}=0.7$).
Figure 4: Comparison of AdamW and GPA on the Llama-8B code generation model using 100 billion tokens.
@@ 4.3 Vision Transformer Model Training
To validate our method on a computer vision task, we train a ViT-S/16 model from timm on ImageNet with data augmentations from the repository. We train this under two batch size settings: (1) a small-batch setup with a batch size of 4,096 for 300 epochs; and (2) a large-batch setup with a batch size of 16,384 for 300 epochs. We tuned the methods separately in both settings, using the average over 2 random seeds to select the best hyperparameters, then ran the best-performing selection on 12 random seeds in total. For all methods, we used gradient clipping with norm 1 and a linear learning rate warmup over the first 5 epochs. We annealed the learning rate with cosine decay to $\times 0.001$ of the peak learning rate for all methods except Schedule-Free, which used constant learning rate after the warmup. Our results are that GPA outperforms AdamW by a clear margin in terms of validation accuracy throughout the course of training (see Figures 5). For further details on our hyperparameter tuning, see Appendix 11.
Figure 5: Comparison of AdamW, DiLoCo, and GPA on ImageNet ViT-S/16 from timm with data augmentations using a batch size of 4,096 samples.
@@ 6 Conclusion
GPA generalizes Nesterov momentum by decoupling interpolation constants for gradient computation and model evaluation. On both small and large-scale language models, GPA outperforms single-worker DiLoCo while eliminating its complex two-loop structure, simplifying hyperparameter tuning, and reducing memory overhead.
@@
Finally, GPA’s decoupling of parameters also enables new avenues for distributed training. In DiLoCo, the number of inner steps serves as a coupled hyperparameter for both Step-$K$ Nesterov and local SGD, leading to the undesirable finding that increasing the number of inner steps can improve convergence – contrary to standard local SGD intuition. By introducing a tunable, continuous smoothing parameter that is independent of the number of inner steps in Step-$K$ Nesterov, GPA provides a new foundation for re-designing algorithms for cross-regional training.